4. Assume that the number of car accidents on Saturday has a Poisson dis- tribution with mean 3.7. Assume that the number of car accidents on Sunday has a Poisson distribution with mean 2.3. Assume that the two random variables are independent. What is the distribution of their sum?
Answer #1

Let, X and Y denote car accidents on Saturday and Sunday respectively, so as per the problem,

X ~ Poisson(A)

y ~ Poisson(μ)

where, lambda and mu are the mean of the poisson distribution X and Y are following respectively.

This only holds if X and Y are independent, so we suppose this from now on. We have for k2 0: k-i k! k! i=0 k!

So it is proved that,

X + y ~ Poisson ( λ + μ)

So the sum will of the poisson distribution will follow poisson distribution with mean lambda + mu

Given the problem, lambda = 3.7 and mu = 2.3

So the sum of two independent poisson distribution is another poisson distribution with mean = 3.7 + 2.3 = 6.